the answer to -9(4b + 5) = -21
Answer by fractalier(6550)

Solving
-9(4b + 5) = -21
we distribute and get
-36b - 45 = -21
Now add 45 and get
-36b = 24
Now divide by -36 and get
b = 24/-36 = -2/3